### v1

@@ 5.1.2 Chest X-ray (COVID-19, Pneumonia) Dataset
Figure 4: Examples from the Chest X-ray (Covid, Pneumonia) dataset (mine, )
5.1.3 CT Kidney Dataset
The CT Kidney dataset consists of 12,446 CT scans in four categories: benign, cyst, stone, and tumor. The images are distributed as follows: 5,077 benign, 3,709 cyst, 1,377 stone, and 2,283 tumor cases. Each CT scan provides cross-sectional insights into the internal anatomy, aiding in diagnosis and analysis. Sample images are shown in Figure 5(kidney, )
@@ 5.1.3 CT Kidney Dataset
Figure 5: Examples from the CT Kidney dataset (mine, )
5.1.4 Diabetic Retinopathy Detection Dataset
This dataset includes 88,645 fundus images, divided into two groups: benign and diabetic retinopathy. There are 65,342 benign and 23,303 diabetic retinopathy images. Fundus photography, utilized here, captures detailed views of the optic nerve, assisting in the detection of abnormalities. Sample images are displayed in Figure 6(diabetic, )
@@ 5.1.4 Diabetic Retinopathy Detection Dataset
Figure 6: Examples from the Diabetic Retinopathy Detection dataset (mine, )
5.1.5 Lung Cancer Histopathological Images
With 15,000 histopathological images across benign, squamous cell carcinoma (SCC), and adenocarcinoma (ACC) classes, this dataset provides 5,000 images per category. Sample images are shown in Figure 7.The data facilitates analysis and classification of histopathological images and was also cited in (mine, ).  (lungcolon, )
@@ 5.1.5 Lung Cancer Histopathological Images
Figure 7: Examples from the Lung Cancer Histopathological Images dataset. (mine, )
@@
In our study, we utilized multiple datasets to thoroughly assess model stability and performance under diverse conditions. The primary focus was on the CT Kidney Dataset, chosen for its consistency in delivering stable, reliable results. Other datasets, such as the Chest X-ray and Diabetic Retinopathy Detection datasets, provided complementary perspectives, allowing us to evaluate model generalizability across different medical imaging types. The CT Kidney dataset’s robustness was especially beneficial, enabling us to prioritize security technique evaluation over model-specific tuning.
6 Experimental Setup for Federated Learning
@@
In this section, we present the experimental setup used to evaluate the performance and security of various cryptographic techniques in a federated learning context on medical image datasets.
6.1 CloudLab Environment
@@
All experiments were conducted on CloudLab (cloudlab, ), a cloud computing testbed designed for systems research. Our setup utilized eleven physical machines in a shared-nothing cluster, connected via Gigabit Ethernet (10 Gbps) to form a local area network. Each machine was equipped with two Intel E5-2683 v3 (14-core) CPUs at 2.00 GHz, 256 GB RAM, and two 1 TB hard drives. Each core had two hardware threads, providing ample processing power for federated learning tasks.
6.2 Federated Learning Framework: Flower
@@
Flower (flower, ) is an open-source federated learning framework, compatible with popular machine learning libraries such as PyTorch (pytorch, ), TensorFlow (tensorflow, ), and MXNet (mxnet, ). In our experiments, we used TensorFlow to train the models and assess accuracy. The Flower framework provided flexible APIs and allowed us to implement various federated optimization strategies. One of the CloudLab machines was designated as the Flower server, while the remaining ten served as Flower clients.
6.3 Federated Learning Setup
@@
Each dataset was divided into 10 equal-sized partitions and distributed across the Flower clients. The centralized federated learning setup consisted of 10 clients and a single server, simulating a typical federated learning framework. Each model was trained over 10 rounds, with each round consisting of 20 epochs. A validation set comprising 10% of the training data was used to monitor model performance. The final model accuracy was evaluated on the test set after completing all rounds.
This controlled environment allowed us to systematically examine the performance and security of different cryptographic methods on decentralized data. Only CPUs were utilized in the setup, creating a realistic resource-constrained environment for federated learning.
6.4 Rationale for Excluding Decentralized Federated Learning
In this study, we opted for a centralized federated learning setup instead of a decentralized approach. A primary consideration was the lack of mutual trust among clients. In decentralized federated learning, clients typically exchange data or model updates directly with one another, creating a network without a central coordinator. However, in our case, clients are unwilling to share their data directly with other participants due to confidentiality concerns and the sensitivity of the information involved.
Centralized federated learning provides a more controlled environment where a central server coordinates the aggregation of model updates, allowing clients to keep their data strictly local while still benefiting from collaborative training. This centralized coordination ensures that data privacy is maintained between clients, as each client only communicates with the trusted server, avoiding direct data exchange with other clients. This setup aligns better with the privacy requirements of our application, where maintaining strict data confidentiality is paramount.
6.5 Deep Learning Models
Deep learning has revolutionized disease diagnosis using medical image datasets, with various convolutional neural networks (CNNs) demonstrating success since 2014. In this study, we utilized five popular models—ResNet-50, DenseNet121, EfficientNetB0, MobileNet V2 for broader evaluation.
6.5.1 ResNet-50
@@
ResNet-50 mitigates the vanishing gradient problem using skip connections, enabling effective training of deep networks and achieving high accuracy in complex tasks.
6.5.2 DenseNet121
@@
DenseNet121 connects each layer to all previous layers, enhancing feature reuse and propagation for efficient, accurate performance in resource-constrained setups.
6.5.3 EfficientNetB0
@@
EfficientNetB0 optimizes accuracy and efficiency by balancing network width, depth, and resolution, making it ideal for computationally constrained scenarios.
6.5.4 MobileNet V2
@@
MobileNet V2 reduces computational demands with lightweight convolutions, maintaining high accuracy for applications on mobile and embedded devices.
- Non Enc: The model is trained without any encryption.
- Full Enc: All data and model weights are fully encrypted using homomorphic encryption.
- Partly Enc: A partly encryption approach is used, encrypting only a portion of the data while scrambling the rest for security.
The overhead columns represent the additional time required for encryption relative to the non-encrypted baseline.
@@
Full Overhead captures the additional time incurred by full encryption, while
Partly Overhead shows the time overhead introduced by partly encryption.
Table 1: Model Parameters and Data Sizes (Fully Encrypted (Full Enc) and Partly Encrypted (Partly Enc))
Model Params (M) Unenc Size (MB) Full Enc (GB) Partly Enc (GB)
EffNetB0 5.3 21.2 2.71 0.281
DenseNet121 7.0 28.0 3.58 0.487
MobileNetV2 3.4 13.6 1.79 0.156
ResNet-50 25.6 102.4 12.4 1.64
Table 1 presents a comparison of model parameters and data sizes for both fully and partly encrypted models. The ”Params” column lists the number of model parameters in millions, while the ”Unenc Size” and ”Enc Size” columns show the data size for unencrypted, fully encrypted, and partly encrypted models. Partly encryption significantly reduces the size of encrypted data compared to full encryption, as illustrated by EfficientNetB0, where the fully encrypted size is 2.71GB, and the partly encrypted size is only 0.281GB.
Table 2: CIFAR-10 Performance Comparison (Time in Minutes)
Model Non Enc Full Enc Partly Enc Full Overhead Partly Overhead
EffNetB0 8m 111m 21m 103m 13m
DenseNet 11m 219m 35m 208m 24m
MobileNet 7m 68m 15m 61m 8m
ResNet-50 16m 530m 68m 514m 52m
Table 2 highlights training times for different models on the CIFAR-10 dataset. Fully encrypted models exhibit significantly longer training times than partly encrypted models due to the overhead of encrypting the entire dataset. For example, DenseNet121 requires 219 minutes to train when fully encrypted, compared to only 35 minutes with partly encryption, illustrating the time-saving benefits of selective encryption without sacrificing data security.
Table 3: Performance on CT Kidney Dataset
Model Non Enc Full Enc Partly Enc Full Overhead Partly Overhead
EffNetB0 112m 228m 127m 116m 15m
DenseNet121 301m 514m 326m 213m 25m
MobileNetV2 91m 163m 100m 72m 9m
ResNet-50 351m 861m 399m 510m 48m
Table 3 shows training times for models on the CT Kidney dataset. The overheads are consistent with the number of model parameters, with heavier models incurring more time overhead. The training overheads for full and partly encryption across models are largely in line with what is expected based on the size and complexity of the models.
Table 4: Performance on Lung Dataset
Model Non Enc Full Enc Partly Enc Full Overhead Partly Overhead
EffNetB0 241m 355m 254m 114m 13m
DenseNet121 287m 506m 312m 219m 24m
MobileNetV2 225m 467m 231m 242m 6m
ResNet-50 319m 834m 372m 515m 53m
Table 4 focuses on the Lung dataset, where training times for encrypted and non-encrypted models are compared. As seen with the CIFAR-10 and CT Kidney datasets, larger models such as ResNet-50 incur the most substantial encryption overhead, while smaller models like MobileNetV2 and EffNetB0 see lighter impacts. The results once again illustrate that encryption overhead correlates directly with the number of parameters and the model’s complexity.
Table 5: Performance on diabetic retinophaty Dataset
Model Non Enc Full Enc Partly Enc Full Overhead Partly Overhead
EffNetB0 578m 698m 591m 120m 13m
DenseNet121 1298m 1542m 1351m 244m 53m
MobileNetV2 546m 610m 555m 64m 9m
ResNet-50 622m 1140m 673m 518m 51m
Table 5 displays the performance on the diabetic retinophaty dataset, which follows the same trend as the previous datasets. Models with more parameters, like ResNet-50, experience higher encryption overheads. The consistency of overhead across different datasets suggests that encryption overheads scale predictably with model size and parameter count, as expected.
6.6 Graphical Comparisons and Analysis
Figure 8: Encryption Overhead Comparison Between Different Models
Figure 8 compares the encryption overheads between fully encrypted and partly encrypted models, revealing a significant reduction in overhead with partly encryption. Partly encryption consistently reduces overhead by approximately 90%, providing a highly efficient method for securing models while still maintaining robust security. This balance between drastically reduced encryption time and maintaining data protection makes partly encryption a compelling alternative to full encryption, especially for larger models, where the time savings are particularly pronounced.
@@ 6.6 Graphical Comparisons and Analysis
Figure 9 illustrates the communication costs associated with different encryption levels for various models, including EfficientNetB0, DenseNet121, MobileNetV2, ResNet-50. The X-axis represents the percentage of encryption applied to the model weights, ranging from 10% to 100% (encryption). The Y-axis displays the corresponding data size in gigabytes (GB), reflecting the communication cost for each model.
The graph shows that as the encryption percentage increases, so does the communication cost. For instance, EfficientNetB0 starts at an unencrypted size of 21.2 MB and grows to 2.71 GB at 100% encryption. Partly encryption mitigates this increase, with significantly lower communication costs at lower encryption levels, such as 0.281 GB for partly encrypted EfficientNetB0.
Across all datasets, fully encrypted models exhibit higher communication costs and computational overhead. Partly encryption, however, reduces both metrics considerably, making it a more practical option for real-time applications where security and performance need to be balanced.
@@
Figure 10 demonstrates that model accuracy remains robust even as the encryption percentage decreases from 90% to 10%. The minimal fluctuations (less than 1%) in accuracy suggest that the encryption technique maintains performance while offering strong security. Any slight variations can be attributed to noise or minor differences in the training process. This experiment was conducted in an offline environment, rather than a federated setup, to isolate and clearly show the effects of encryption on model accuracy. dataset used in this accuracy representation is lung cancer dataset.
Summary: Accuracy remains largely unaffected by partly encryption, maintaining a stable performance across different models. The minimal decrease in accuracy demonstrates that partly encryption provides a viable alternative to full encryption, offering security without compromising model effectiveness.
Figure 11: Training Time and Encrypted Size Comparison for Fully and Partly Encrypted Models For CT Kidney dataset
Figure 11 compares the training times and encrypted sizes of both fully and partly encrypted models for CT Kidney dataset. The graph visually illustrates the significant reduction in encryption overhead when applying partly encryption techniques compared to full encryption.For larger models like ResNet-50, fully encrypted models show a sharp increase in both training time and encrypted size.
FAS delivers substantial improvements in performance efficiency, reducing training times and encrypted sizes across a variety of datasets. It emerges as a viable strategy for federated learning where computational resources are limited, yet security remains a top priority.
6.7 Effects on Security
Before conducting our main security experiments, we implemented several security tests to evaluate the robustness of our encryption methods. These tests included integrity checks, timing attack tests, noise and precision evaluations, and simulations of chosen-ciphertext and chosen-plaintext attacks. Additionally, differential attack tests and statistical distribution checks were performed to ensure that encryption did not compromise data integrity or privacy. Based on these initial tests, we found that 10% partly encryption successfully passed the same security tests as 100% full encryption, indicating that partly encryption maintains a high level of security while reducing computational demands.
Table 6: Full vs Partly Encryption: Security Test Results
@@ 6.7 Effects on Security
Table 6 presents the results of comprehensive security tests comparing partly encryption with scrambling to full encryption. Both encryption schemes underwent a variety of cryptographic attacks and precision tests to evaluate their robustness. The results indicate that partly encryption offers comparable security outcomes to full encryption across most criteria, without significant compromises in protection.
6.7.1 Evaluating Privacy with MSSIM and VIFP
@@
Model inversion attacks threaten federated learning by reconstructing training data from model updates. To evaluate privacy robustness, we utilize the Mean Structural Similarity Index (MSSIM) and Visual Information Fidelity in the Pixel domain (VIFP).
@@
MSSIM quantifies structural similarity between original and reconstructed images, with low scores indicating effective privacy measures. Similarly, VIFP assesses visual quality based on human perceptual models, detecting fine-grained artifacts and the impact of techniques like differential privacy noise and bitwise scrambling. Both metrics are widely recognized for their ability to evaluate privacy in adversarial scenarios.
@@
Our results show that selective encryption combined with scrambling and noise significantly degrades reconstruction quality, with MSSIM stabilizing at 52% for 20% encryption and VIFP confirming similar trends across multiple encryption levels. These findings highlight the effectiveness of our approach and the utility of MSSIM and VIFP as reliable privacy evaluation metrics.
Following these tests, we further evaluated the effectiveness of partly encryption against model inversion attacks. In these attacks, we attempted to reconstruct images from the encrypted models and compared the reconstructed images to the originals using VIFP (vifp, ) and MSSIM (mssim, ) metrics.
Despite using only 10% encryption, the reconstructed images displayed significant distortion. For instance, at 10% encryption, the MSSIM score was 58%, indicating substantial deviation from the original image. As the encryption percentage increased, the score stabilized at around 52%, further demonstrating that partly encryption provides strong protection against inversion attacks, even with reduced encryption levels.
@@
Figure 12: MSSIM Scores across Encryption Percentages for Various Models
@@
Figure 12 illustrates the MSSIM scores for various models across different levels of partly encryption. The displayed scores were obtained by averaging results from multiple experimental runs to account for randomness in selective encryption, scrambling, and noise. As encryption increases from 10%, there is a sharp decline in MSSIM scores, with most models stabilizing around 20% encryption. These results suggest that even low levels of partly encryption can achieve security comparable to full encryption, with significantly reduced computational demands. Notably, at 100% encryption, the resistance against inversion attacks is very close to the resistance achieved at 10% and similar to that at 20% encryption, highlighting the effectiveness of selective encryption even at lower levels.
Figure 13: VIFP Scores across Encryption Percentages for Various Models
Figure 13 shows VIFP scores across different encryption percentages for various models. As with MSSIM, VIFP scores drop sharply at 10% encryption and stabilize at around 20%. The displayed scores were repeated across multiple experimental runs to ensure robustness and mitigate the impact of randomness introduced by scrambling and noise. Importantly, the resistance to model inversion attacks at 100% encryption** is very similar to that achieved at 10% and 20% encryption levels. This stability indicates that partly encryption offers effective protection against model inversion attacks, even at lower encryption levels.
6.8 Cross-testing with related work
6.8.1 Comparison with MASKCRYPT (10506637, )
MASKCRYPT is a selective encryption method designed to minimize the overhead of fully encrypting model updates by applying a gradient-guided encryption mask. Instead of pre-training, MASKCRYPT calculates gradient values during each round, determining which model weights are critical for privacy protection and selectively encrypting those. This reduces communication overhead by up to 4.15× compared to full model encryption and decreases membership inference attack accuracy to 49.2% (ma2020secure, ). However, MASKCRYPT’s gradient mask is not always optimized for selecting the most sensitive gradients, which can result in lower initial security scores.
Figure 14: MSSIM and VIFP scores over rounds : MASKCRYPT
Figure 14 depicts the VIFP scores for MASKCRYPT over five rounds of federated learning. The MASKCRYPT model was recreated by us to evaluate its performance. While the results may not fully match the original models values, they offer a reliable insight into MASKCRYPT’s behavior. As observed in the graph, MASKCRYPT requires multiple rounds to achieve an acceptable encryption stability level, with a noticeable reduction in VIFP scores between rounds 1 and 3, eventually stabilizing around round 5.
This gradual improvement contrasts with our proposed FAS method, which stabilizes at a much earlier stage. This difference in behavior further highlights the efficiency of our approach in achieving secure and stable encryption without requiring as many iterations.

### v2

@@ V-A2 Chest X-ray (COVID-19, Pneumonia) Dataset
Figure 4: Examples from the Chest X-ray (Covid, Pneumonia) dataset [20]
V-A3 CT Kidney Dataset
Comprising 12,446 CT scans categorized as benign (5,077), cyst (3,709), stone (1,377), and tumor (2,283), this dataset aids in kidney disease diagnosis [9].
@@ V-A3 CT Kidney Dataset
Figure 5: Examples from the CT Kidney dataset [20]
V-A4 Diabetic Retinopathy Dataset
This dataset comprises 88,645 fundus images, categorized as benign (65,342) or diabetic retinopathy (23,303). Fundus imaging assists in detecting optic nerve abnormalities [6].
@@ V-A4 Diabetic Retinopathy Dataset
Figure 6: Examples from the Diabetic Retinopathy Detection dataset [20]
V-A5 Lung Cancer Histopathological Images
This dataset contains 15,000 histopathological images, equally distributed across benign, squamous cell carcinoma (SCC), and adenocarcinoma (ACC) classes [8].
@@ V-A5 Lung Cancer Histopathological Images
Figure 7: Examples from the Lung Cancer Histopathological Images dataset. [20]
@@
Our evaluation primarily focused on the CT Kidney dataset, chosen for its stable, reliable results. The Chest X-ray and Diabetic Retinopathy datasets provided additional perspectives to assess model generalizability across different medical imaging contexts. The CT Kidney dataset’s robustness enabled prioritization of security technique evaluation over model-specific tuning.
VI Experimental Setup for Federated Learning
@@
This section details the experimental setup for evaluating various cryptographic techniques in a federated learning context on medical image datasets.
VI-A CloudLab Environment
@@
Experiments were conducted on CloudLab [10], a cloud computing testbed for systems research. Our setup utilized 11 physical machines in a shared-nothing cluster, connected via 10 Gbps Ethernet. Each machine was equipped with dual Intel E5-2683 v3 CPUs (14-core, 2.00 GHz), 256 GB RAM, and two 1 TB hard drives, providing substantial computational resources for federated learning tasks.
VI-B Federated Learning Framework: Flower
@@
Flower [11], an open-source federated learning framework, was used for model training and evaluation, supporting TensorFlow [12], PyTorch [13], and MXNet [14]. One CloudLab machine served as the Flower server, while the remaining 10 functioned as Flower clients.
VI-C Federated Learning Setup
@@
Each dataset was divided into 10 equal partitions, distributed across the Flower clients. The centralized federated learning setup comprised 10 clients and a server. Each model was trained over 10 rounds, with 20 epochs per round. A validation set (10% of training data) monitored performance, and final accuracy was evaluated on a test set. All training used CPUs to simulate resource-constrained environments.
VI-D Rationale for Centralized Federated Learning
Decentralized federated learning was not used due to a lack of mutual trust among clients. A centralized setup ensures data privacy, as clients communicate only with the server, avoiding direct data exchange and aligning with confidentiality requirements.
VI-E Deep Learning Models
We evaluated four deep learning models:
VI-E1 ResNet-50
@@
ResNet-50 mitigates the vanishing gradient problem with skip connections, facilitating deep network training.
VI-E2 DenseNet121
@@
DenseNet121 enhances feature reuse by connecting each layer to all previous layers, improving efficiency in resource-constrained settings.
VI-E3 EfficientNetB0
@@
EfficientNetB0 optimally scales network width, depth, and resolution, balancing accuracy and efficiency.
VI-E4 MobileNet V2
@@
MobileNet V2 employs lightweight convolutions to reduce computational costs while maintaining accuracy, making it suitable for mobile and embedded applications.
VI-F Encryption Configurations
We compared three encryption configurations:
  • Non-Enc: No encryption applied.
  • Full-Enc: All model weights encrypted using homomorphic encryption.
  • Partly-Enc: Selective encryption applied to a subset of data, with scrambling for security.
@@ VI-F Encryption Configurations
Full Overhead represents additional training time due to full encryption, while Partly Overhead captures the impact of partial encryption.
TABLE I: Training Time (Minutes) and Overheads Across Datasets
Model Dataset Full Enc (min) Partly Enc (min)
EffNetB0 CIFAR-10 111 21
CT Kidney 228 127
Lung 355 254
COVID 698 591
Diabetic Retinopathy 1298 1351
DenseNet121 CIFAR-10 219 35
CT Kidney 514 326
Lung 506 312
COVID 1542 1351
Diabetic Retinopathy 1542 1351
MobileNetV2 CIFAR-10 68 15
CT Kidney 163 100
Lung 467 231
COVID 610 555
Diabetic Retinopathy 610 555
ResNet-50 CIFAR-10 530 68
CT Kidney 861 399
Lung 834 372
COVID 1140 673
Diabetic Retinopathy 1140 673
Table I presents training times (minutes) across datasets. Fully encrypted models require significantly more training time, with ResNet-50 showing the highest overhead. In contrast, partly encrypted models reduce overhead substantially while maintaining security.
@@
Figure 8 shows communication costs across encryption levels, revealing higher costs for fully encrypted models, whereas partly encrypted models optimize performance.
@@
Figure 9 indicates that accuracy remains stable across encryption percentages, confirming the effectiveness of partly encryption in preserving model performance.
Figure 10: Training Time and Encrypted Size Comparison for Fully and Partly Encrypted Models (CT Kidney dataset)
Figure 10 highlights significant reductions in encryption overhead when using partly encryption, making FAS a viable option for resource-limited federated learning scenarios.
VI-G Effects on Security
Before conducting our main security experiments, we implemented several security tests to evaluate the robustness of our encryption methods. These tests included integrity checks, timing attack tests, noise and precision evaluations, and simulations of chosen-ciphertext and chosen-plaintext attacks. Additional differential attack tests and statistical distribution checks were performed to verify that encryption preserved data integrity and privacy. Our findings indicate that 10% partly encryption successfully passed the same security tests as 100% full encryption, demonstrating strong security with reduced computational overhead.
TABLE II: Full vs Partly Encryption: Security Test Results
@@ VI-G Effects on Security
Table II compares partly encryption with full encryption across a series of security evaluations. The results confirm that partly encryption provides comparable protection, reinforcing its viability as a lightweight yet effective security mechanism.
VI-G1 Evaluating Privacy with MSSIM and VIFP
@@
Model inversion attacks pose a threat to federated learning by reconstructing training data from model updates. To assess privacy resilience, we use Mean Structural Similarity Index (MSSIM) and Visual Information Fidelity in the Pixel domain (VIFP).
@@
MSSIM quantifies structural similarity between original and reconstructed images, with lower scores indicating stronger privacy. VIFP measures visual quality based on perceptual models, highlighting the impact of noise and scrambling techniques. Both metrics are widely recognized for privacy evaluation in adversarial scenarios.
@@
Our results show that selective encryption with scrambling and noise significantly degrades reconstruction quality. MSSIM stabilizes at 52% for 20% encryption, and VIFP follows a similar trend across multiple encryption levels. This confirms that our approach effectively mitigates inversion attacks while optimizing computational efficiency.
Further tests demonstrated that even at 10% encryption, reconstructed images showed significant distortion. The MSSIM score at 10% encryption was 58%, highlighting a substantial deviation from the original images. As encryption increased, the score stabilized at around 52%, reinforcing the effectiveness of partial encryption in maintaining privacy.
@@
Figure 11: MSSIM and VIFP Scores across Encryption Percentages for Various Models
@@
Figure 11 illustrates MSSIM and VIFP scores across encryption percentages for various models. Results were averaged across multiple runs to account for variations introduced by selective encryption, scrambling, and noise. The sharp decline in scores at 10% encryption, with stabilization around 20%, suggests that lower encryption levels can provide security equivalent to full encryption while significantly reducing computational costs. Notably, resistance to inversion attacks at 100% encryption is nearly identical to that at 10% and 20%, highlighting the robustness of selective encryption.
VI-H Cross-testing with Related Work
VI-H1 Comparison with MASKCRYPT
MASKCRYPT [19] applies a gradient-guided encryption mask to selectively encrypt critical model weights, reducing communication overhead by up to 4.15× compared to full model encryption. However, its gradient mask is not always optimized for selecting the most sensitive gradients, leading to lower initial security scores. MASKCRYPT requires multiple rounds to stabilize encryption, as reflected in VIFP scores, which only stabilize by round 5. This indicates an adaptation period where security is suboptimal.